OpenStudy (anonymous):

What am I doing wrong here? (5x^2 - 5) / (x - 1) = 5x(x - 1) / (x - 1) = 5x The answer should be 5(1 + 1)

OpenStudy (anonymous):

answer should be 10??? O_O

OpenStudy (anonymous):

5(x+1) Sorry, typo!

OpenStudy (callisto):

(5x^2 - 5) / (x - 1) = 5(x + 1) / (x - 1) ________ The underlined part is wrong. 5x^2-5 Take out the common factor 5 => 5 (x^2 -1) Factor (x^2-1) using the identity \(a^2-b^2 = (a+b)(a-b)\) Can you try it again?

OpenStudy (callisto):

Oops sorry! The underlined part should be 5(x-1)

OpenStudy (anonymous):

Aha, got it! Thanks

OpenStudy (callisto):

Welcome :)